OpenStudy (astrophysics):

Find the values of x for which the series converges. Find the sum of the series for those values of x.

OpenStudy (astrophysics):

\[\sum_{n=1}^{\infty} (x+2)^n\]

OpenStudy (astrophysics):

@ganeshie8

OpenStudy (dan815):

x<3 is my first guess

OpenStudy (astrophysics):

You love your x?

Parth (parthkohli):

An infinite series converges when the last term approaches zero.

Parth (parthkohli):

lol, was this a trick question?

OpenStudy (astrophysics):

Nope

Parth (parthkohli):

x < 3? x = 2 doesn't seem to work.

OpenStudy (astrophysics):

What's the common ratio?

Parth (parthkohli):

Well, I'd say that\[-1 < x+2 < 1\]

OpenStudy (astrophysics):

That's right

OpenStudy (dan815):

thats not how u solve it, its a geometric series with some shift

OpenStudy (astrophysics):

that's the interval of convergence

OpenStudy (dan815):

x+2 is the ratio

OpenStudy (dan815):

let y=x+2 sum of n from 1 to inf of y^n

OpenStudy (astrophysics):

Yeah, so |r|=|x+2|<1 is when it converges

Parth (parthkohli):

Great. For the series to converge, you want the common ratio's abs value to be less than 1.

Parth (parthkohli):

Yeah.

Parth (parthkohli):

And\[S = \dfrac{a}{1-r} = \dfrac{x+2}{1 - x - 2}\]

OpenStudy (astrophysics):

Yup

OpenStudy (dan815):